OpenStudy (anonymous):

Describe how to transform the graph of f into the graph of g. f(x) = square root x-2 g(x) = square root x+6

OpenStudy (debbieg):

f(x-h) will translate the graph of f(x), h units to the right if h>0, and |h| units to the left if h<0. So write \(\sqrt{x+6}\) as \(\sqrt{x-2-h}\) ... what does h=? That will tell you what you need to do with the graph of f to get the graph of g.

OpenStudy (debbieg):

Another way to do it is to think of the x-intercepts.... what x gives you f(x)=0? What x gives you g(x)=0? Where are those x's in relation to one another?

OpenStudy (debbieg):

(that is more the intuitive way; the first way is more the rigorous mathematical way :)

OpenStudy (anonymous):

Shift the graph of f left 8 units.

OpenStudy (anonymous):

is tht the answer

OpenStudy (debbieg):

Yes, that's it.